OpenStudy (anonymous):

combine the following rational expressions: (1)/(8x^3-1) - (1)/ (4x^2-1)

OpenStudy (anonymous):

have you first tried facoting the denominators of both expresions

OpenStudy (anonymous):

yes but you cant can you? thers no x with the 1

OpenStudy (anonymous):

well, we have to notice something here: recall these two formulas: Difference of square: (a^2-b^2)=(a-b)(a+b) Difference of cubes: a^3-b^3=(a-b)(a^2+ab+b^2)

OpenStudy (anonymous):

now lets dissect your two denominators: First, we have (8x^3-1) can i write this as: (2x)^3-(1)^3

OpenStudy (anonymous):

if i can (which i can) then it we can factor this using the formula i wrote up there, for the difference of cubes

OpenStudy (anonymous):

The same goes for your other denominator ,expecet that one is a difference oif squares(i wrote the formula up there for this one two) I can write (4x^2-1) as (2x)^2-(1)^2

OpenStudy (anonymous):

now once you factor these two you will see what you need to multiply each fraction by inorder to subtract them

OpenStudy (anonymous):

ok thanks!